### v1

# Omnigenous umbilic stellarators

Abstract
To better understand the dependence of the magnetic field structure in the plasma edge due to plasma boundary shape, we define and develop umbilic stellarators. These equilibria are characterized by a single continuous high-curvature edge on the plasma boundary that goes around multiple times toroidally before meeting itself. We develop a technique that allows us to simultaneously optimize the plasma boundary along with a curve lying on the boundary on which we impose a high curvature while imposing omnigenity — a property of the magnetic field that ensures trapped particle confinement throughout the plasma volume. After generating omnigenous umbilic stellarators, we design coil sets for some of these equilibria and calculate and understand the field line structure in the edge. Finally, we propose an experiment to modify an existing tokamak to a stellarator using this technique and explore a potential way to convert a limited tokamak into a diverted stellarator.
1 Introduction
Stellarators (Spitzer Jr, 1958) are toroidal nuclear fusion devices that are used to study magnetic confinement fusion. Unlike the axisymmetric tokamaks, stellarators possess discrete toroidal symmetry. The lack of continuous toroidal symmetry offers considerably greater design freedom, at the expense of trapped particle confinement, a quality guaranteed in axisymmetry through Noether’s theorem (Noether, 1918). Therefore, we seek stellarators with trapped particle confinement, a property known as omnigenity.
Omnigenous magnetic fields ensure zero bounce-averaged radial drift (Hall & McNamara, 1975; Cary & Shasharina, 1997), of which both axisymmetry and quasisymmetry are subsets (Boozer, 1983; Nührenberg & Zille, 1988; Dudt et al., 2024). Recent advances in stellarator optimization have generated omnigenous configurations with excellent neoclassical confinement levels(Landreman & Paul, 2022; Goodman et al., 2023). However, these configurations are typically only calculated up to the Last Closed Flux Surface (LCFS). To develop a stellarator Fusion Power Plant (FPP), it is essential to also predict and optimize the magnetic field structure outside the LCFS (Parra Diaz et al., 2024). This is necessary to separate the high-temperature plasma from the Plasma Facing Components (PFCs) while efficiently disposing of fusion by-products like helium ash and impurities such as heavier ions.
The structure outside the plasma boundary depends on the boundary shape, so changing the boundary shape is a potential way to obtain a desirable fieldline structure in the edge. Indeed, divertors (Burnett et al., 1958) have been greatly successful in tokamaks and offer significantly improved performance over limiters (Wagner et al., 1982). Depending on the boundary shape and coil positions, one can design various divertor configurations in tokamaks such as the single-null, double-null, snowflake (Ryutov, 2007), and super-X (Valanju et al., 2009).These configurations utilize poloidal and divertor coils to create and shift an X point — a region in tokamaks where the poloidal field disappears.
Stellarators have utilized three distinct divertor configurations (Bader et al., 2018a): helical, resonant, and non-resonant divertors. Helical divertors were used in heliotron/torsatron concepts such as the Heliac device (Harris et al., 1985) and LHD (Ohyabu et al., 1992). These divertors work by creating a stochastic region with edge surface layers and X-points, which the magnetic field lines cross before reaching the scrape-off layer(SOL). The coil design and magnetic configuration in the edge surface layers is complex and difficult to predict (Ohyabu et al., 1994). Moreover, applicability of this concept to omnigenous stellarators with non-helical symmetry still remains uncertain (Bader et al., 2018a), posing challenges for their design and optimization.
Resonant divertors, used in W7-AS (Grigull et al., 2001; Feng et al., 2002) and W7-X (Pedersen et al., 2019), work by leveraging island formation and placing divertors between the X and O points of the islands. The island structure depends on the pitch of the field line in the boundary, whereas the connection length depends directly on the number of X points and inversely on the magnetic field line pitch and the magnetic shear (Feng et al., 2011) making them useful for low magnetic shear stellarators.
Non-resonant divertors use a more complicated design to create a stochastic field in the edge, but similar to resonant divertors, they can be used to spread the load and improve plasma exhaust rejection. These divertors are also more resilient to changes in the magnetic field structure (Bader et al., 2017), but the magnetic field structure in the chaotic field region is complicated to calculate and requires further understanding. Recent work (Boozer & Punjabi, 2018; Punjabi & Boozer, 2022; Garcia et al., 2023; Davies et al., 2025) has further explored the possibility of using non-resonant divertors in stellarators.
Motivated by the need to understand how boundary shaping affects fieldline structure in the edge and divertor design, in this paper, we study the umbilic stellarator (US) as a method of achieving control over the edge magnetic field topology, and generate umbilic stellarators for vacuum and finite-$\beta$ cases111where $\beta=2\mu_{0}p/B^{2}$ is the ratio of the plasma to the magnetic pressure. Umbilic stellarators are characterized by regions of high curvature on the boundary and how these high-curvature regions are positioned. We develop a technique that allows us to obtain umbilic stellarators by simultaneously optimizing a curve and the plasma boundary, and imposing a high curvature along this curve which creates a ridge along the plasma boundary. Additionally, we ensure that all equilibria have reasonable omnigenity, by ensuring a low neoclassical transport. Finally, we propose an experiment to modify an existing tokamak to a stellarator using this technique and build it using umbilic coils. We show how an umbilic coil, different from helical coils, can create an umbilic edge in finite-pressure equilibria, acting as an alternate way to create a ridge on the boundary.
In §2, we explain the ideal MHD equilibrium. In §3, we introduce the umbilic boundary shape, how we approximate them with DESC, and how we solve the ideal MHD equations throughout the volume. In §4, we present a set of optimized vacuum($\beta=0$) and finite-$\beta$ stellarators and analyze their properties such as omnigenity. In §5, we take the Columbia HBT-EP plasma shape, modify it into an umbilic stellarator, and design electromagnetic coils for it. In §6, we conclude our work and discuss how it can be extended.
2 Ideal MHD equilibrium
In this section, we briefly explain how we define and calculate an ideal MHD equilibrium in a stellarator. We then explain how to locally vary the gradients of that equilibrium.
A divergence-free magnetic field $\boldsymbol{B}$ can be written in the Clebsch form (D’haeseleer et al., 2012)
$\boldsymbol{B}=\nabla\psi\times\nabla\alpha,$ (1)
We will focus on solutions whose magnetic field lines lie on closed nested toroidal surfaces, known as flux surfaces. We label these surfaces using the enclosed toroidal flux $\psi$. On each flux surface, the lines of constant $\alpha$ coincide with the magnetic field lines. Thus, $\alpha$ is known as the field line label. We define $\alpha=\theta_{\rm{PEST}}-\iota\zeta$, where $\theta_{\rm{PEST}}$ is the $\mathrm{PEST}$ straight field line angle, $\zeta$ is the cylindrical toroidal angle, and
$\iota=\frac{\boldsymbol{B}\cdot\boldsymbol{\nabla}\theta_{\rm{PEST}}}{\boldsymbol{B}\cdot\boldsymbol{\nabla}\zeta},$ (2)
is the pitch of the magnetic field lines on a flux surface, known as the rotational transform. Using the Clebsch form of the magnetic field, we solve the ideal MHD force balance equation
$\boldsymbol{j}\times\boldsymbol{B}=\boldsymbol{\nabla}p,$ (3)
where the plasma current $\boldsymbol{j}=(\boldsymbol{\nabla}\times\boldsymbol{B})/\mu_{0}$ from Ampere’s law, $p$ is the plasma pressure and $\mu_{0}$ is the vacuum magnetic permeability. Unlike an axisymmetric case, for stellarators we have to solve (3) as an optimization problem. We achieve this with the DESC (Dudt & Kolemen, 2020; Panici et al., 2023; Conlin et al., 2023; Dudt et al., 2023) stellarator optimization suite through minimization of $F\equiv\boldsymbol{j}\times\boldsymbol{B}-\boldsymbol{\nabla}p$. DESC can simultaneously solve an equilibrium while optimizing for multiple objectives such as MHD stability, quasisymmetry and more, without re-solving the equilibrium force balance equation (3) for each optimization variable. To reduce the computational cost further, we also use the fact that stellarators possess a discrete toroidal symmetry such that every stellarator comprises several identical sections. The number of identical sections is denoted by the field period $n_{\rm{FP}}$. Therefore, the force balance problem is only solved in a single field period.
In the following sections, we utilize DESC to optimize for omnigenous umbilic stellarators.
3 Umbilic stellarators
Umbilic stellarators draw inspiration from an umbilic bracelet (or torus) (Zeeman, 1976), a 3D shape with a cross-section that is a deltoid curve, a three-cusped hypocycloid which rotates by $\Delta\theta=2\pi/3$ poloidally after a single toroidal turn $\Delta\zeta=2\pi$. The sharp edge lies exactly on the cusp. Therefore, the sharp edge has to go around three times toroidally before meeting itself. A stellarator generated using an umbilic boundary is potentially advantageous for divertor designs, offering a continuous divertor option with a long connection length. For a smooth magnetic field, the sharp edge of such a boundary would be impossible for a magnetic field line. Therefore, the fieldline pitch on the boundary matches the umbilic topology That is, $\iota=n_{\rm{FP}}m/n=1/3$, where $m,n\in\mathbb{Z}$ corresponding to the poloidal and toroidal motion of the fieldline and $n_{\rm{FP}}$, the field period, is the measure of discrete symmetry of the stellarator. For a general umbilic shape, $m/n$ can attain rational values besides $1/3$. Further details are provided in this section.
3.1 Defining an umbilic surface and edge
The initial shape of a stellarator symmetric umbilic-torus-like surface $\partial V$ is parameterized as a rotating polygon-like with convex sides
$\begin{split}R_{\mathrm{US}}&=R_{0}+2\varrho\cos\left(\frac{\pi}{2n}\right)\cos\left(\frac{\theta+\frac{\pi}{n}(2\lfloor\frac{n\theta}{2\pi}\rfloor+1)}{2}+\frac{mn_{\rm{FP}}\zeta+r_{1}\sin(n_{\rm{FP}}\zeta)+r_{4}\sin(2n_{\rm{FP}}\zeta)}{n}\right)\\ &-\varrho\cos\left(\frac{\pi}{n}(2\Big{\lfloor}\frac{n\theta}{2\pi}\Big{\rfloor}+1)+\frac{mn_{\rm{FP}}\zeta+r_{1}\sin(n_{\rm{FP}}\zeta)+r_{4}\sin(2n_{\rm{FP}}\zeta)}{n}\right)+r_{2}\cos(n_{\rm{FP}}\zeta)\end{split}$ (4)
$\begin{split}Z_{\mathrm{US}}&=2\varrho\cos\left(\frac{\pi}{2n}\right)\sin\left(\frac{\theta+\frac{\pi}{n}(2\lfloor\frac{n\theta}{2\pi}\rfloor+1)}{2}+\frac{mn_{\rm{FP}}\zeta+r_{5}\sin(n_{\rm{FP}}\zeta)}{n}\right)\\ &-\varrho\sin\left(\frac{\pi}{n}(2\Big{\lfloor}\frac{n\theta}{2\pi}\Big{\rfloor}+1)+\frac{mn_{\rm{FP}}\zeta+r_{5}\sin(n_{\rm{FP}}\zeta)}{n}\right)+r_{3}\sin(n_{\rm{FP}}\zeta)\end{split}$ (5)
where the major radius $R_{0}=1$ throughout this paper, unless stated otherwise, $\varrho$ is the minor radius of the umbilic-torus-like surface, $\theta$ is the poloidal angle used in DESC, $\zeta$ is the cylindrical toroidal angle, and $\lfloor$..$\rfloor$ … ⌋ is the floor function, defined such that
$\lfloor x\rfloor=\begin{cases}x,&\text{if }x\in\mathbb{Z}\\ N-1,&\text{if }N-1<x<N\text{ and }x\notin\mathbb{Z},\ N\in\mathbb{Z}\end{cases}$
In the surface parametrization, we also define $n$ as the umbilic factor, which defines the number of sides in the analytical parametrization, $m$ is an integer that determines the poloidal motion of the sharp umbilic edge, and $n_{\rm{FP}}$ is the field period. The periodicity of the edge is $n_{\rm{FP}}(m/n)$ — the edge closes on itself after $n/\gcd(n,n_{\rm{FP}})$ toroidal turns, where $\gcd$ is the greatest common divisor. The parameters $r_{1},r_{4},r_{5}$ define the convexity/concavity of a cross-section and how it changes toroidally, $r_{4}$ determines the magnetic axis torsion and $r_{2}$ is another parameter that determines the shape of the axis. The shape of such a boundary, the naming convention we shall use, and periodicity of the edge are illustrated in figure 1.
(a) US432 full surface
(b) US432 single field period
[…]
The floor function creates the sharp umbilic edge, while the rest of the parameters grant us enough freedom that we can start with a boundary shape that has good quasisymmetry and allows us to create an umbilic edge with any low-order rational periodicity $n_{\rm{FP}}(m/n)$. It is also stellarator symmetric, i.e., $R_{\mathrm{US}}(\theta,\zeta)=R_{\mathrm{US}}(-\theta,-\zeta),Z_{\mathrm{US}}(\theta,\zeta)=-Z_{\mathrm{US}}(-\theta,-\zeta)$.
We also define a separate grid on the boundary $\rho=1$, the umbilic edge, on which a sharp curvature is imposed, defined by values $(\hat{\theta},\zeta)$ so that
$\hat{\theta}=\frac{mn_{\rm{FP}}\zeta+\sum_{k}^{N_{\mathrm{u}}}a_{k}\sin(n_{\rm{FP}}\zeta)}{n},\quad\gcd(m,n)=1,$ (6)
For the surface parametrization, $\theta=0$ corresponds to the umbilic edge — initial $\hat{\theta}$ is calculated along this edge but it can change freely during optimization as the optimizer varies the coefficients $a_{k}$. Note that $\theta$ is the DESC poloidal angle throughout this paper, unless stated otherwise.
However, creating and solving with a perfectly sharp edge is difficult with a fully spectral solver. Therefore, in the next section, we will explain how we approximate umbilic boundary shapes with the DESC code.
3.2 Approximating umbilic shapes with DESC
DESC uses a Fourier-Zernike spectral representation to solve the force balance equation in the $(\rho,\theta,\zeta)$ coordinate system, where $\rho=\sqrt{\psi/\psi_{b}},\,\theta=\theta_{\rm{PEST}}-\Lambda(\rho,\theta,\zeta),\,\zeta=\phi$, the cylindrical toroidal angle. The problem is defined in a cylindrical coordinate system $(R,\zeta,Z)$ by decomposing it into Fourier-Zernike spectral bases as shown below (See (Panici et al., 2023) for more details)
$R(\rho,\theta,\zeta)=\sum_{m=-M,n=-N,l=0}^{M,N,L}R_{lmn}\mathcal{Z}_{l}^{m}(\rho,\theta)\hat{F}^{n}(\zeta)$ (7)
$Z(\rho,\theta,\zeta)=\sum_{m=-M,n=-N,l=0}^{M,N,L}Z_{lmn}\mathcal{Z}_{l}^{m}(\rho,\theta)\hat{F}^{n}(\zeta)$ (8)
where $l,m$, and $n$ are the radial, poloidal, and toroidal mode numbers whereas $L,M$, and $N$ define the largest values of $l,m$, and $n$, respectively and $\mathcal{Z},\hat{F}$ are Zernike and Fourier basis functions. These parameters define the resolution of a DESC equilibrium.
To solve the ideal MHD equation inside the umbilic shape, we have to first fit the Fourier-Zernike basis at $\rho=1$ to the umbilic boundary, which requires solving a set of linear equations.
$\centering\left[\begin{array}[]{cc}\sum_{m=-M,n=-N,l=0}^{M,N,L}R_{lmn}\mathcal{Z}_{l}^{m}(\rho=1,\theta)\hat{F}^{n}(\zeta)\\[5.69054pt] \sum_{m=-M,n=-N,l=0}^{M,N,L}Z_{lmn}\mathcal{Z}_{l}^{m}(\rho=1,\theta)\hat{F}^{n}(\zeta)\end{array}\right]=\left[\begin{array}[]{cc}R_{\mathrm{US}}(\theta,\zeta)\\[5.69054pt] Z_{\mathrm{US}}(\theta,\zeta)\end{array}\right],\@add@centering$ (13)
to find $R_{lmn},Z_{lmn}$ so that the boundary shape matches with the umbilic parametrization. Similarly, to obtain $a_{k}$, we parameterize the umbilic edge using $R_{\mathrm{curve}}=R(\rho=1,\theta=0,\zeta),Z_{\mathrm{curve}}=Z(\rho=1,\theta=0,\phi)$, and perform an inverse Fourier transform to obtain $a_{k}$. Note that upon fitting a spectral basis, the umbilic edge is smoothed out. After representing the umbilic shape in Fourier-Zernike basis, we solve the ideal MHD equation inside the volume. Figure 2 illustrates this process of fitting and solving force balance in a US131 boundary shape.
Figure 2: Contour plot of the flux surface average force balance error $|F|$ plotted on the cross section $\zeta=0$ of a vacuum DESC equilibrium created using a US131 boundary with $\rho=0.1,r_{1}=0.133,r_{5}=0.1$. The actual boundary (in solid black) is obtained after fitting a Fourier-Zernike series to the true umbilic boundary(in dashed blue).The resolution of the DESC equilibrium is $L=M=N=14$ and the resolution of the umbilic curve $N_{\mathrm{u}}=12$.
The sharpness of the umbilic edge increases as we increase the resolution of the fit surface and the agreement between DESC and the umbilic parametrization improves. Alternatively, one may use a double Chebyshev (Fortunato, 2024) instead of a double Fourier representation to improve the spectral fit, local bases methods such as finite-element to represent and solve the equilibrium, or use a non-linear transformation to map the circular torus in Fourier-Zernike basis to an umbilic shape, similar to the transformation used for circular tokamaks by Webster & Gimblett (2009). However, using a new basis and performing equilibrium optimization is beyond the scope of this work.
In the following section, we discuss how beginning with the parametrization of the umbilic surface, we can optimize for omnigenous umbilic stellarators.
3.3 Optimizing an umbilic equilibrium with DESC
As we noted in the previous section and figure 2, after fitting a Fourier-Zernike basis to the umbilic boundary, it becomes smooth because approximating a sharp boundary with a double-Fourier representation is usually difficult. Therefore, we have to modify the optimization process for an umbilic equilibrium to maintain a high curvature, instead of a perfectly sharp umbilic edge. Additionally, for these shapes to be practically feasible fusion reactors, we also need good omnigenity. Therefore, we use the omnigenity and effective ripple objectives to improve the omnigenity throughout the volume.
After fitting to a Fourier-Zernike basis, the high-curvature edge, in general, does not align with the field line. In the unoptimized shape defined in (4) and (5), the $\iota$ is typically too small because of weak shaping. Therefore, we have to optimize separately for a boundary rotational transform to be close $n_{\rm{FP}}m/n$. Also, to ensure that the umbilic edge curve matches with the field line labeled $\alpha=0$, we enforce a strong negative curvature $\kappa_{2,\rho}$ along the umbilic edge defined by (6). The definitions of the principal curvatures are given in Appendix A. The overall objective function is
$\mathcal{F}_{\mathrm{stage-one}}=w_{A}\,f_{\mathrm{aspect}}^{2}+w_{\iota}\,f_{\rm{iota}}^{2}+w_{O}\,f_{\mathrm{om}}^{2}+w_{U}\,f_{\mathrm{umbilic}}^{2}+w_{R}\,f_{\mathrm{ripple}}^{2},$ (14)
where $f_{x}$ are various objectives on the right side for aspect ratio, boundary rotational transform, omnigenity, high curvature along the umbilic edge, and effective ripple and $w_{A},w_{\iota},w_{O},w_{U},w_{R}$ are weights used with each objective function. The exact definitions of these objectives are provided in appendix B. At each iteration of a DESC optimization, $\mathcal{F}_{\mathrm{stage-one}}$ is minimized while satisfying (3). Mathematically,
$\min\mathcal{F}_{\mathrm{stage-one}}(\hat{\boldsymbol{p}}),\qquad\textrm{s.t.}\quad\boldsymbol{\nabla}\left(\mu_{0}p+\frac{B^{2}}{2}\right)-\boldsymbol{B}\cdot\boldsymbol{\nabla}\boldsymbol{B}=0,\quad\psi_{\mathrm{b}}=\psi_{\mathrm{b}0},$ (15)
where $\psi_{\mathrm{b}0}$ is the user-provided enclosed toroidal flux by the boundary and $\hat{\boldsymbol{p}}\in\{R_{\mathrm{b},mn},Z_{\mathrm{b},mn},a_{k}\}$ comprises parameters that determine the boundary shape and the position of the umbilic curve on the boundary. With the objective $\mathcal{F}_{\mathrm{stage-one}}$ defined, we run DESC on a single NVIDIA A100 GPU. A single optimization takes less than $90$ minutes.
4 Omnigenous umbilic stellarators
In this section, we will develop two omnigenous umbilic equilibria using the process explained in the previous sections. The first equilibrium will be a single field period vacuum stellarator with $n=3,m=1$ umbilic topology for which we also design coils, and the second equilibrium will be a two field period finite-$\beta$ stellarator with $n=5,m=2$ umbilic topology.
4.1 US131($n_{\rm{FP}}=1,n=3,m=1$) omnigenous vacuum stellarator
We begin by calculating an omnigenous vacuum equilibrium with a single field period $n_{\rm{FP}}=1$, $n=3,m=1$. To do this, we start with the US131 boundary shape, using the parametrization from section 3.1 with $\hat{\rho}=0.1\rm{m},r_{1}=0.133,r_{5}=0.1$ and the rest of the parameters $r_{j}=0$. The toroidal flux enclosed by the initial equilibrium $\psi_{\rm{b}}=\pi\hat{\rho}^{2}(\rm{T}m^{2})$. This shape is then approximated by DESC as explained in section 3.2 and optimized as described in section 3.3 with $w_{O}=0$  [^2] By setting $w_{O}=0$ and $w_{R}\neq 0$, we seek omnigenous configurations without constraining the magnetic field to a specific helicity. This improves the quality of the optimum. This could also be a reason we are more likely to get piecewise-omnigenous configurations. to obtain an omnigenous equilibrium with an aspect ratio $A=10.5$. The results are presented in figures 3 and 4.
(a) Cross section comparison
(b) Rotational transform
(c) Alignedness$(\rho=1.0)$
Figure 3: US131 optimization results showing a significant increase in magnetic axis torsion in figure (a) from the initial (dotted) to the optimized (solid) boundary. This leads to an increase in the rotational transform, as shown in figure (b) such that $\iota(\rho=1)$ approaches $1/3$. The shaping also leads the fieldlines to align with the umbilic edge in the optimized equilibrium as shown in figure (c).
To ensure that the high-curvature umbilic edge is aligned with the magnetic field lines, the boundary undergoes deformation, which results in an increased rotational transform. The optimized design minimized the objective $\mathcal{F}_{\mathrm{stage-one}}$ while ensuring that the normalized second principal curvature is the lowest ($\kappa_{2,\rho}<-35\rm{m}^{-1}$) along the umbilic edge. This causes the magnetic field line to stay close to the umbilic edge throughout the range $\zeta\in[0,6\pi]$ as shown in figure 3(c). In the limit where the edge is perfectly sharp, the field line will coincide with the umbilic edge.
The optimized US case is able to achieve an effective ripple $\epsilon_{\mathrm{eff}}$ of less than $2\%$ ($\epsilon_{\mathrm{eff}}^{3/2}<0.0028$) for most of the volume $(\rho<0.86)$. This ensures that neoclassical transport in the $1/\nu$ low-collisionality regime will not deteriorate the device performance. An interesting observation is that we can achieve a low ripple transport despite the lack of a specific helicity of the $B$ contours in the Boozer plot figure $4\mathit{(b)}$. This is an instance of a piecewise omnigenous (Velasco et al., 2024) field that is far from quasisymmetry. We prove this by analyzing the normalized second adiabatic invariant $\tilde{\mathcal{J}}_{\parallel}$ in Appendix C.
(a) Initial B ($\rho=1$)
(b) Optimized B ($\rho=1$)
(c) Neoclassical transport
Figure 4: Output from the vacuum US131 optimization showing Boozer plots on the boundary and the effective ripple. Figure (b) does not have a specific helicity but still achieves a low ripple transport.
For a vacuum problem, obtaining the right fieldline pitch ($\iota=n_{\rm{FP}}m/n$) on the boundary and ensuring force balance over a perfectly sharp (unapproximated) boundary makes this problem difficult to solve compared to a finite-$\beta$ equilibrium where one can specify the field line pitch throughout the volume. Therefore, we have compromised by creating a high-curvature edge instead of a perfectly sharp edge ($\kappa_{2,\rho}\rightarrow\infty$). If one were to use an alternate representation to represent a perfectly sharp boundary and solve this problem exactly, we would observe the rotational transform almost identical to the initial equilibrium in figure 4(c) sharply rising to $\iota=m\,n_{\rm{FP}}/n$ at the edge.The behavior of high-shear regions near sharp points would be similar to what is observed in tokamaks.
4.2 Modular coil design for the US131 equilibrium
In this section, we will design modular coils for the US131 case using the DESC optimizer and elucidate the various difficulties faced when designing modular coils for strongly-shaped umbilic stellarators. We create a coilset of $32$ (stellarator-symmetric with $16$ unique coils) coils using the Python implementation of REGCOIL (Landreman, 2017; Panici et al., 2025) in DESC. Starting with this initial coilset, we then perform stage-two optimization, where we fix the equilibrium boundary and allow the shape of the modular coils to change while minimizing the objective
$\mathcal{F}_{\rm{stage-two}}=(B_{\rm{out}}^{2}-(B_{\rm{in}}^{2}+2\mu_{0}p))^{2}+(\boldsymbol{B}_{\rm{out}}\cdot\boldsymbol{\hat{n}})^{2}$ (16)
on the plasma boundary. Note that in stage-two optimization, we ignore the surface current $\boldsymbol{K}=(\boldsymbol{B}_{\rm{out}}-\boldsymbol{B}_{\rm{in}})\times\boldsymbol{\hat{n}}$. Therefore, to verify our solution we calculate the field line trajectory due to the coils and create a Poincaré plot. The original fixed-boundary solution, its comparison with the Poincaré cross-section, and the optimized coilset are presented in figure 5.
Figure 5: Coils and $\boldsymbol{B}_{\mathrm{out}}\cdot\hat{\boldsymbol{n}}$ error plotted on the plasma boundary. The $\mathrm{max}|(\boldsymbol{B}_{\mathrm{out}}\cdot\hat{\boldsymbol{n}})/\boldsymbol{B}_{\mathrm{out}}|\leq 1\%$. Only one group of unique coils are plotted. On the upper right we plot the Poincaré cross-section at $\zeta=0$ along with the fixed-boundary solution(in red). The gray curves mark a favorable position to place a divertor. In the lower right part we plot the $\boldsymbol{B}_{\mathrm{out}}\cdot\hat{\boldsymbol{n}}$ error on a 2D plane along with the umbilic edge(dashed purple).
Modular coil design becomes complicated for umbilic equilibria as seen in the figure. The coils assume high curvature to reduce the normal field error along the umbilic edge. Increasing the number of coils improves the agreement of the fixed boundary solution with the Poincaré cross-section and reduces the max coil curvature but also reduces the coil-coil distance. However, the number of modular coils needed for these umbilic equilibria already exceeds the number of coils used for a single field period for a typical stellarator, which is almost always less than ten.
Another important point is that even though the field lines do not exit in the regions of the highest curvature, the Poincaré traces stay aligned with the plasma boundary and flare near the umbilic edge, which still makes the umbilic edge a favorable location to put divertors or limiters.
[…]
To ensure practical coil design for umbilic stellarators, a better technique would be to perform single-stage optimization, which simultaneously changes the plasma boundary and coil shapes, a union of $\mathcal{F}_{\mathrm{stage-one}}$ and $\mathcal{F}_{\mathrm{stage-two}}$, accounting for the coil complexity while maintaining omnigenity inside the plasma volume. Another possibility is to use umbilic coils, helical coils with a fractional helicity($m/n$) that create cusp-like regions, leading to a high curvature umbilic edge. We do a similar optimization to modify a tokamak in section 5, where we add an umbilic coil without changing the shape of the toroidal field coils to create an umbilic edge. In the next section, we shall use the same optimization process to create a finite-$\beta$, omnigenous umbilic stellarator.
4.3 US252 $(n_{\rm{FP}}=2,n=5,m=2)$ omnigenous finite-$\beta$ stellarator
In this section, we will generate a two field period $n=5,m=2$ finite-$\beta$ stellarator. We start with the umbilic parametrization defined in subsection 3.1 with $\varrho=0.1,r_{1}=0.133,r_{5}=0.1$ and use a pressure profile $p=5\times 10^{3}(1-\rho^{2})\mathrm{Pa}$, a rotational transform $\iota=0.755+0.043\rho^{2}$, and an enclosed toroidal flux $\psi_{\rm{b}}=\pi\hat{\rho}^{2}=0.00314\,\rm{T}m^{2}$. This creates an equilibrium with a $\beta=0.5\%$.
A finite-$\beta$ stellarator provides more freedom compared to a vacuum case because the rotational transform can come from a combination of boundary shaping and plasma current. Therefore, we can achieve a higher rotational transform and use the additional freedom to achieve a much lower ripple transport compared to the vacuum case presented in the previous section. The optimization uses the same objective function defined in (14) with the pressure and iota profiles fixed and $w_{\iota}=0$. These results are presented in figure 6.
(a) Initial $|B|(\rho=1.0)$
(b) Optimized $|B|(\rho=1.0)$
(c) Neoclassical transport
Figure 6: Optimization output showing the initial and final magnetic field strengh Boozer plots in figures (a) and (b), respectively and the effective ripple in figure (c). The optimized configuration is another instance of piecewise omnigenity
We can see in figure 6, the ridges are only clearly visible in the initial equilibrium, which is close to quasisymmetry. The magnetic field of the optimized omnigenous equilibrium does not have any visible ridges even though the curvature along the umbilic ridge increases after optimization. This indicates that the appearance of ridges may only be a feature of quasisymmetric stellarators.
This equilibrium is optimized for poloidal omnigenity but, due to the umbilic curvature penalty, limits the poloidal omnigenity to the low-field region. The Boozer plot in figure 6(b) looks qualitatively similar to the piecewise-omnigenous configuration of the W7-X stellarator. The piecewise omnigenity of this equilibrium is demonstrated further in Appendix C. We also plot the cross-section of the equilibrium along with the second principal curvature in figure 7.
(a) Optimized cross-sections
(b) Curvature on boundary
(c) $\iota$ fraction from shaping
Figure 7: Cross-section and three-dimensional shape of the optimized US252 equilibrium. The aspect ratio of the optimized equilibrium is $A=6.5$. In figure (c), we see that around half of the rotational transform of the optimized equilibrium comes from shaping (Hirshman & Hogan, 1986) reducing the dependence on external current drive.
We can see that the optimized equilibrium deforms and rotates as we move toroidally, while maintaining high curvature along the umbilic curve. The variation in the curvature is broad, indicating that imposing high curvature at specific locations may be more suitable than others. This can be readily accomplished by selecting sections of the umbilic curve where high curvature is applied and could be a potential way to place potential locations for non-resonant divertors (Bader et al., 2018b).
5 Umbilic coil design for a finite-$\beta$ equilibrium
As we have seen in section 4.1, creating modular coils for vacuum umbilic equilibria is a difficult task due to the sharp curvature regions. Therefore, in this section, we present an alternate strategy that uses umbilic coils to create the right boundary shape for a finite-$\beta$ equilibrium. Instead of modifying the modular coils, we can use an umbilic coil, a low-current helical coil which can be used to create an umbilic edge. In the following sections, we will discuss two cases, categorized by the direction of the current in the umbilic coils with respect to the toroidal plasma current.
5.1 Umbilic coil current opposite to the plasma current
Umbilic coils differ from helical coils by creating a ridge that may serve as a potential location for an X-point. These coils are similar to tokamak divertor coils, albeit without axisymmetry. To demonstrate their applicability, we modify a $n=1$ MHD-stable equilibrium from the High-Beta Torus-Extended Pulse (HBT-EP) experiment ($R_{0}=0.94m,A=7.5$) by adding an $m=1,n=1$ ridge to it, breaking its axisymmetry, and then design an umbilic coil for it without modifying the toroidal or vertical field coils. The specific details of the equilibrium used are given in Appendix D.
To accomplish this, we start by using $20$ Toroidal Field (TF)  [^4] In the HBT-EP experiment the TF coils have been wound multiple times and packed into a square Aluminum frame with finite thickness of over $0.15\rm{m}$ whereas we assume filamentary TF coils in our calculations, ignoring the effect of finite thickness. The radius of the TF filamentary coils is $0.65\rm{m}$, which is the average of the outer and inner dimensions of a real HBT-EP TF coil. coils and $4$ Vertical Field (VF) poloidal coils as done in the original HBT-EP experiment (Gates, 1993; Sankar et al., 1993) and add a $m=1,n=1$ umbilic coil outside the plasma boundary at a distance of $1.5$ times the minor radius. Similar to section 4.2, we perform second stage optimization to find the right shape of the umbilic coil. To ensure that the coils from stage-two optimization give expected results, we calculate the free-boundary solution and present it in figure 8.
(a) Mod. HBT-EP full coilset
(b) $\kappa_{2,\rho}$ umbilic coil
(c) X-section & umbilic coil
Figure 8: Output from the stage-two optimization and free-boundary solve. Figure (a) shows the full coilset of the modified HBT-EP experiment with the magnetic field strength on the boundary whereas figure (b) shows the umbilic coil and the curvature $\kappa_{2,\rho}$ on the boundary. Figure (c) shows the cross section of the LCFS from the free-boundary DESC solver along with the umbilic coil at four different toroidal angles and the directions in which they carry current relative to the plasma.
The effect of the umbilic coil can be decomposed as a rigid shift of the whole equilibrium and a small deformation of the boundary. Since the rigid shift (Sengupta et al., 2024) and the size of the ridge is small compared to the minor radius, it does not compromise omnigenity (as seen later in figure 11(a)). For a perfectly sharp umbilic edge, the magnetic field line must coincide with the edge, but even when the ridge lacks perfect sharpness, the magnetic field line stays close to it due to its high curvature.
[…]
To understand the structure of the magnetic field near the edge, we use the FIELDLINES  (Lazerson et al., 2016) code from the STELLOPT (Lazerson et al., 2020) optimizer.
(a) Poincaré section ($\zeta=0$)
(b) Poincaré section ($\zeta=\pi/2$)
(c) Poincaré section ($\zeta=\pi$)
Figure 9: Poincaré section at different toroidal angles compared with free-boundary DESC solution(red). The red cross marks the position of the umbilic coil.
We observe the presence of a stochastic layer outside the plasma boundary, similar to heliotron/torsatron configurations. Since the plasma and the umbilic coil are carrying currents in opposite directions, the poloidal field cannot vanish between them, which is why we do not observe an X-point. Therefore, the umbilic coil can create and vary the curvature of the ridge without creating an varying the current in the umbilic coil without creating an X point. This also affects the fraction of the rotational transform generated by the umbilic coils. These details are provided in figure 6.
(a) $I_{\mathrm{umbilic}}=4.1kA$
(b) $I_{\mathrm{umbilic}}=6.1kA$
(c) $\iota$ fraction from shaping
Figure 10: Free-boundary solution at $\zeta=0$ for different values of the umbilic coil current $I_{\mathrm{umbilic}}$ shown in (a)-(b) and the fraction of the rotational transform generated by the umbilic coil current [^6] Since all the coils except the umbilic coil are axisymmetric and circular, any departure from a circular plasma shape is a consequence of the umbilic coil. Furthermore, for a perfectly circular equilibrium ($\iota_{\mathrm{shaping}}=0$), so all the rotational transform on the boundary must come from non-circular shaping. Therefore, we conclude that any finite $\iota_{\mathrm{shaping}}$ a consequence of the umbilic coil.in (c).
By controlling the current in the umbilic coil, we can shape the plasma without moving the limiter. Also, the ability to vary $\iota$ would help to stabilize various MHD instabilities and the $m=1,n=1$ tearing mode as done in the RFX-mod experiment (Piron et al., 2016) or experiments like the Madison Spherical Torus, where the edge rotational transform close to one can be achieved (Hurst et al., 2022). The J-TEXT (Liang et al., 2019) tokamak has recently proposed a similar idea (Li et al., 2024) of using umbilic coils to stabilize the edge of a tokamak.
Upon breaking the axisymmetry of the magnetic field, one must be careful as increasing the current $I_{\mathrm{umbilic}}$ for the same coil shape would degrade omnigenity and hence neoclassical transport. To understand the effect of umbilic coils on omnigenity, we present Boozer plots for the three values of current in figure 11.
Figure 11: Magnetic field strength in Boozer coordinate on the boundary $(\rho=1.0)$ for different values of the umbilic coil current and the fieldline $\alpha=0$. As the curvature of the ridge increases, so does the localized distortion of the $|\boldsymbol{B}|$ contours close to the fieldline.
We can see that the omnigenity degrades as we increase $I_{\mathrm{umbilic}}$ but the neoclassical losses are still acceptable $\epsilon_{\mathrm{eff}}^{3/2}<0.002$. To control instabilities and transport due to MHD modes, we may need to generate a higher $\iota_{\mathrm{shaping}}$ while ensuring omnigenity. One way to explore multiple such coil and equilibrium configurations would be to perform a single-stage optimization, which simultaneously changes the plasma boundary and coil shapes. This is beyond the scope of the current study.
5.2 Umbilic coil current in the same direction as the plasma current
In a manner similar to the previous section, we create a setup where we modify the HBT-EP experiment and add a $m=n=1$ ridge to the boundary. However, we do this while ensuring that the toroidal component of the current in the umbilic coil $I_{\mathrm{umbilic}}=-6.1kA$ is in the same direction as the plasma current $I=-35.2kA$. The coilset and free-boundary solution are presented in figure 12.
(a) Mod. HBT-EP full coilset
(b) $\kappa_{2,\rho}$ umbilic coil
(c) X-section & umbilic coil
Figure 12: Coilset and free-boundary solution for the case with the coils are closer to the plasma and the current in the umbilic coil points in the same direction as the plasma current and since no stage-two optimization was performed the second-stage objective $\mathcal{F}_{\mathrm{stage-two}}$ has a large value.
Due to the proximity of the coil to the plasma, the maximum $\lvert\boldsymbol{B}\cdot\hat{\boldsymbol{n}}\rvert$ and pressure balance errors were larger than the acceptable normalized value (of $1\%$), implying the boundary shape will differ from the one shown in figure 12.
[…]
Since the umbilic coil can reduce the poloidal field significantly, we hypothesize that it should be possible to optimize for the coil position and plasma boundary shape so that there is a region between the plasma and the coil that behaves like an X point. This could potentially convert the initially limited HBT-EP into a diverted configuration.
6 Conclusions and future work
This paper explored the physical properties of umbilic stellarators. In section 3, we defined the umbilic parametrization, approximated it using DESC and solved the ideal MHD equation inside the enclosed volume. In section 4, we developed vacuum and finite-$\beta$ stellarators by simultaneously optimizing the umbilic edge and the boundary shape using the DESC optimization suite. We designed coils for the vacuum case, and by tracing the field lines in the edge, showed how the high-curvature edges could be a favorable location to place divertors. We found that the optimized omnigenous umbilic designs do not possess a specific helicity regardless of the low-ripple transport, which could be a way to generate piecewise omnigenous equilibria. As a final demonstration, in section 5, we propose an experiment to modify the Columbia HBT-EP experiment to an umbilic design by adding a $m=1,n=1$ helical coil. Using the DESC free boundary solver, we demonstrate that the umbilic coils are able to cause the originally circular plasma to bulge and form a high-curvature ridge while maintaining omnigenity. By changing the direction of the current in the umbilic coil, we may be able to get an X-point, thereby converting the limited HBT-EP tokamak into a diverted stellarator.
The technique of simultaneously optimizing a curve and a surface to control the magnitude and position of a sharp edge along a surface extends beyond umbilic stellarators. One can use this method to place and design ridges or regions of high curvature which may be useful for non-resonant divertors (Bader et al., 2018b), as well as these ridges creating preferential pathways for the field lines to exit the plasma (Bader et al., 2017). For an island divertor concept, the rotational transform on the boundary can be fine-tuned using umbilic coils, which could be used to reduce the sensitivity of X-points in the island divertor structure due to changes in the plasma boundary rotational transform. This idea is similar to that of divertor coils in tokamaks and would lead to a predictable strike point pattern.
[…]
Acknowledgements One of the authors, R. G. enjoyed discussions with M. Zarnstroff, J. Levesque, J. L. Velasco, B. Jang, and gratefully acknowledges the help from S. Lazerson installing the FIELDLINES code. This work is funded through the SciDAC program by the US Department of Energy, Office of Fusion Energy Science, and Office of Advanced Scientific Computing Research under contract No. DE-AC02-09CH11466, the Hidden Symmetries grant from the Simons Foundation/SFARI (560651), and Microsoft Azure Cloud Computing mini-grant awarded by the Center for Statistics and Machine Learning (CSML), Princeton University. This research also used the Della cluster at Princeton University.
Data availability statement The data and analysis files are freely available in Gaur (2025).
Appendix A Calculating principal curvatures
The second principal curvature at a given point on a surface is defined as a
$\centering\kappa_{2,\rho}=\min\left\{k:\det\left[\begin{array}[]{cc}L_{\mathrm{sff},\rho}-kE&M_{\mathrm{sff},\rho}-kF\\ M_{\mathrm{sff},\rho}-kF&N_{\mathrm{sff},\rho}-kG\end{array}\right]=0\right\}\@add@centering$ (17)
where
$\begin{split}E&=\frac{\partial\boldsymbol{x}}{\partial\zeta}\cdot\frac{\partial\boldsymbol{x}}{\partial\zeta},\quad F=\frac{\partial\boldsymbol{x}}{\partial\zeta}\cdot\frac{\partial\boldsymbol{x}}{\partial\theta},\quad G=\frac{\partial\boldsymbol{x}}{\partial\theta}\cdot\frac{\partial\boldsymbol{x}}{\partial\theta},\\ L_{\mathrm{sff},\rho}&=\frac{\partial^{2}\boldsymbol{x}}{\partial\theta^{2}}\cdot\frac{\boldsymbol{\nabla}\rho}{|\boldsymbol{\nabla}\rho|},\quad M_{\mathrm{sff},\rho}=\frac{\partial^{2}\boldsymbol{x}}{\partial\theta\partial\zeta}\cdot\frac{\boldsymbol{\nabla}\rho}{|\boldsymbol{\nabla}\rho|},\quad N_{\mathrm{sff},\rho}=\frac{\partial^{2}\boldsymbol{x}}{\partial\zeta^{2}}\cdot\frac{\boldsymbol{\nabla}\rho}{|\boldsymbol{\nabla}\rho|}\end{split}$ (18)
are the metric coefficients corresponding to the first and second fundamental forms. The user-specified values $\kappa_{2,\mathrm{bound1}},\kappa_{2,\mathrm{bound2}}$ define the limits on the curvature. Using this objective, we impose a maximum and minimum value on $\kappa_{2,\rho}$ that helps us impose a sharp curvature along the umbilic edge.
Appendix B Form of various objectives
All the objectives with bounds are defined as
$f_{\mathrm{obj}}=\mathrm{ReLU}(\max(\rm{obj}-\rm{obj}_{\mathrm{bound1}},\rm{obj}-\rm{obj}_{\mathrm{bound2}})).$ (19)
In this paper, such a term is used to penalize the aspect ratio, rotational transform, and the second principal curvature $\kappa_{2,\rho}$ along the umbilic edge.
For the effective ripple objective explained in Unalmis et al. (Unalmis et al., 2024) and the omnigenity objective implemented by Dudt et al. (Dudt et al., 2024), we do not use any bounds and try to reduce the magnitude of these quantities as much as possible. The exact definitions are given in the respective papers.
Appendix C Piecewise omnigenity in umbilic configurations
[…]
(a) US131 $(\rho=0.5)$
(b) US252 $(\rho=0.75)$
(c) $n_{\rm{FP}}\!=\!1,\!A\!=\!5$ OP$(\rho\!=\!1)$
Figure 13: Plots of the second adiabatic invariant $\mathcal{J}_{\parallel}(\alpha,1/\lambda)$ and $|B|$ contours for three equilibria with $\epsilon_{\mathrm{eff}}^{3/2}\approx 0.001$. For a close-to-perfect omnigenous equilibrium $\partial_{\alpha}\mathcal{J}_{\parallel}$ close to zero for most values of $1/\lambda$, as seen in (c). However, for (a) and (b) there are multiple distinct values of $\mathcal{J}_{\parallel}$, discontinuous in $\alpha$, a necessary conditions for piecewise omnigenity.
For a perfect, non-piecewise omnigenous stellarator $\partial_{\alpha}(\mathcal{J}_{\parallel})=0$, which means that $\mathcal{J}_{\parallel}$ is continuous and constant in $\alpha$ for almost every $1/\lambda$. Moreover, the distinct $\mathcal{J}_{\parallel}$ can be written as a linear sum such that the total $\mathcal{J}_{\parallel}$ for a value of the inverse pitch vanishes. However, for the umbilic configurations in this paper, especially the US131 vacuum stellarator, despite a low effective ripple (good omnigenity), we clearly see that $\mathcal{J}_{\parallel}$ is discontinuous in $\alpha$ for most of the inverse pitch values and $\tilde{\mathcal{J}}_{\parallel,\mathrm{I}}\approx\tilde{\mathcal{J}}_{\parallel,\mathrm{II}}\approx 0.5\tilde{\mathcal{J}}_{\parallel,\mathrm{III}}$. This supports our claim that the umbilic stellarator US131 is piecewise omnigenous, especially at large values of the inverse pitch. US252 seems to be piecewise omnigenous but the structure of $\tilde{\mathcal{J}_{\parallel}}$ is more complicated than the one discussed by Velasco et al. (2024).
Appendix D Details of the HBT-EP equilibrium
The original HBT-EP equilibrium is a shifted-circle equilibrium with a major radius $R_{0}=0.94m$, minor radius $a_{0}=0.125m$ (aspect ratio $A=7.52$) with a $\beta=0.5\%$ and a plasma current $I=-35.2kA$. The pressure and rotational transform profiles are given in figure 14. The initial equilibrium was tested to be stable against $n=1$ kink modes using the DCON-2D (Glasser, 2016) stability code.
(a) Pressure
(b) Rotational Transform
Figure 14: Profiles of the initial HBT-EP equilibrium.